1. L1 and L2 are lines in R
3
. A parametric equation for L1 is
(x, y, z) = (1, 0, 2) + t(2, 4, −1); t ∈ R.
L2 is parallel to L1 and contains the point (3, 4, 3). V is the plane that contains both the lines L1 and L2. Find
an equation for V . [7]
2. Prove from first principles that
lim
(x,y)→(2,2)
(x − y) = 0.
[8]
3. Prove that lim
(x,y)→(0,0)
x
2 + y
x − y
does not exist. [8]
4. Consider the R
2 − R function f defined by
f (x, y) = x
2 − 2xy + y
2
.
Let V be the plane that is tangent to the graph of f at the point (x, y, z) = (1, −1, 4).
(a) Find a vector that is perpendicular to the plane V . (4)
(b) Find an equation for the plane V .
1
Expert's answer
2021-10-04T16:23:17-0400

1,

"\overline{a}=(2,4,-1)||V"

"A(1,0,2)\isin L_1"

"B(3,4,3)\isin L_2"

"\overline{AB}=(2,4,1)||V"

"\overline{n}=[\overline{a},\overline{AB}]=\begin{vmatrix}
 i & j&k \\
 2 & 4&-1\\
2&4&1
\end{vmatrix}=8i-4j=(8,-4,0)"

equation for V:

"8(x-1)-4y=0"

"2x-y-2=0"


2.

 Let "\varepsilon" > 0 be given.

triangle inequality:

"|x-y|=|(x-2)-(y-2)|\le|x-2|-|y-2|"

Also:

"|x-2|\le\sqrt{|x-2|^2}\le|(x-2),(y-2)|"

and

"|y-2|\le|(x-2),(y-2)|"

 Therefore:

"|x-y|\le2|(x-2),(y-2)|"

 Set "\delta=\varepsilon/2" , then if

"0<2|(x-2),(y-2)|<\delta"

we have

"|y-2|\le|(x-2),(y-2)|<2\delta=\varepsilon"


3.

 On the curve C1: x=0

"\displaystyle{\lim_{(x,y)\to(0,0)}}\frac{x^2+y}{x-y}=\displaystyle{\lim_{(x,y)\to(0,0)}}\frac{y}{-y}=-1"


On the curve C2: y=0

"\displaystyle{\lim_{(x,y)\to(0,0)}}\frac{x^2+y}{x-y}=\displaystyle{\lim_{(x,y)\to(0,0)}}\frac{x^2}{x}=0"


 As the limits are not equal, "\frac{x^2+y}{x-y}"  cannot have a limit at (0,0).


4.

a)

Let

"g(x,y,z)=x^2-xy-y^2-z"

"gradg" is normal to the graph of f.

"gradg(x,y,z)=(2x-2y,2y-2x,-1)"

Therefore "gradg(1, \u22121, 4) = (4, \u22124, 1)" will be normal to the tangent plane V .


b)

"(x,y,z)(4,-4,-1)=(1,-1,4)(4,-4,-1)"

"4x-4y-z=4"